I need help on find the third angle of a right triangle if one of the angles measures 24 degree

Answer by macston(5194)

One angle must be 90degrees to be a right triangle, one measures 24 degrees, and the sum of the angles in a triangle is 180 degrees.
A+B+C=180 degrees
24 degrees + B + 90 degrees=180 degrees
B + 114 degrees=180 degrees Subtract 114 degrees from each side.
B=66 degrees ANSWER: The third angle measures 66 degrees.